#1. Do you know how to solve for 3m + 2m = 15
#2. Do you know how to solve for 4r + 5r +45
#3. Do you know how to solve for 10p - 7p = 18
#4. Do you know how to solve for 2a + 9a = 88
#5. Do you know how to solve for 5y + 4y - 11 = 16

Answer by rfer(16322)

1)
3m+2m=15
5m=15
m=15/5
m=3
3)
10p-7p=18
3p=18
p=18/3
p=6
5)
5y+4y-11=16
9y=16+11
9y=27
y=27/9
y=3
You try the others